OpenStudy (anonymous):

What is the product of 7 sqrt(10) * 3 sqrt(5)?

OpenStudy (anonymous):

\[7\sqrt{10} \times 3\sqrt{5}\]

OpenStudy (tkhunny):

Hint: \(\sqrt{10} = \sqrt{2}\cdot\sqrt{5}\)

OpenStudy (anonymous):

I know but neither of those are perfect square numbers...

OpenStudy (jdoe0001):

\(\bf 7\sqrt{10}\times 3\sqrt{5}\implies (7\times 10)(\sqrt{10\times5})\)

OpenStudy (jdoe0001):

hmmm...got a typo :(

OpenStudy (anonymous):

It's multiple choice and these are the options a. 21the square root of 5. b. 105the square root of 2. c. 105the square root of 5. d. 21the square root of 2.

OpenStudy (jdoe0001):

\(\bf 7\sqrt{10}\times 3\sqrt{5}\implies (7\times 3)(\sqrt{10\times5})\)

OpenStudy (tkhunny):

Rule #1 - If the first thing that crosses your mind doesn't seem to work, DO NOT give up. Think harder and think of something else. \(\sqrt{10}\cdot\sqrt{5} = \sqrt{2}\cdot\sqrt{5}\cdot\sqrt{5} = 5\sqrt{2}\) You must look around and see whatever else there is.

OpenStudy (anonymous):

Yeah but that one doesn't work either...

OpenStudy (tkhunny):

Really? You have to do ALL the work, not just part of it. \(7\cdot\sqrt{10}\cdot3\cdot\sqrt{5} = 7\cdot3\cdot\sqrt{10}\cdot\sqrt{5} = 21\cdot5\cdot\sqrt{2} = 105\cdot\sqrt{2}\) Seriously, hang in there a little longer. Giving up too soon is a very bad habit.

OpenStudy (shamil98):

7√10 = 7 x (√2 √5) x 3√5 Combine the like terms (7x3) ( √5 x √5) 21 x 5 x √2 = 105√2